OpenStudy (anonymous):

Can somebody find the average rate of change? y = (x-1)^2 -3

OpenStudy (anonymous):

Here is a picture of the graph:

OpenStudy (anonymous):

I need to find the average rate of change from x=-1 to x=1

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you want change in y divided by change in x find y at x=-1, and y at x=1 now you have two (x,y) points that you can use

OpenStudy (anonymous):

I don't understand @phi

OpenStudy (phi):

can you find the y value when x=-1 using y = (x-1)^2 -3 ?

OpenStudy (anonymous):

Y would equal 1

OpenStudy (anonymous):

Y=1

OpenStudy (phi):

ok, so you have (-1,1) as a point on the curve now y when x=+1

OpenStudy (anonymous):

Y=-3?

OpenStudy (phi):

so (1,-3) is another point on the curve now find the slope between those two points. That is the average rate of change between the two points.

OpenStudy (phi):

Here is a picture

OpenStudy (anonymous):

Is the slope 2? The picture is confusing me...

OpenStudy (phi):

you have two points (-1,1) and (1,-3) find the change in y divided by the change in x can you do that ?

OpenStudy (anonymous):

-1,-1/3?

OpenStudy (phi):

using (-1,1) and (1,-3) the change in y is 1 - (-3) = 1 + 3 = 4 the change in x is -1 - 1 = -2 change in y divided by change in x is 4/-2 = -2

OpenStudy (phi):

that means that on average, your curve goes down 2 steps for every step to the right, between those two points.

OpenStudy (anonymous):

Oh, I did (-1,1) / (1,-3)

OpenStudy (anonymous):

I see you're a moderator now

OpenStudy (phi):

If this stuff is unclear, see http://www.khanacademy.org/math/cc-eighth-grade-math/cc-8th-relationships-functions/cc-8th-slope/v/slope-of-a-line-2 and his other videos on how to find the slope between two points.

OpenStudy (anonymous):

Thank you so much Phi